OpenStudy (anonymous):

which fraction is equivalent to 9/16

OpenStudy (anonymous):

there could be many... One that's smaller or larger?

OpenStudy (anonymous):

What are the options?

OpenStudy (paki):

simple... just see... 9/16 = 18/32 = 27/48 = 36/64 = 45/80 = 54/96 = 63/112 = 72/128 = 81/144 = 90/160 = 99/176 = 108/192 = 117/208 = 126/224 = 135/240 = 144/256 = 153/272 = 162/288 = 171/304 = 180/320 and so on......